A roasted turkey is taken from an oven when its temperature has reached 185oF and is placed on a table in a room where the temperature is 75oF.
a) If the temperature of the turkey is 150oF after half an hour, what is its temperature after 40 min?
b) When will the turkey cool to 90oF?

how do you solve this college algebra problems?:)?
i have answers but i don't know how to solve it. could you explain in details please?

Answer by stanbon(57214)

A roasted turkey is taken from an oven when its temperature has reached 185oF and is placed on a table in a room where the temperature is 75oF.
-----
You have 1 point relating time and temperature:
(0,185)

a) If the temperature of the turkey is 150oF after half an hour, what is its temperature after 40 min?
You now have a 2nd point relating time and termperature:
(30,150)
-----
Find the equation relating time and temp:
intercept = 185
slope + (185-150)/(0-30) = -35/30 = -7/6
----
Equation: f(t) = (-7/6)t+185
what is its temperature after 40 min?
f(40) = (-7/6)(40)+185 = 138.33 degrees Farenheit
-------------------------------
b) When will the turkey cool to 90oF?
90 = (-7/6)t+185
(-7/6)t = -95
t = 85.43 minutes
====================
Cheers,
Stan H.
====================